How many grams of sodium hydroxide would you need to produce 75 g of lead (II) hydroxide? 2 NaOH(aq) + Pb(NO3)2(aq) → 2 NaNO3(aq) + Pb(OH)2(s)
1
Expert's answer
2020-05-23T12:26:29-0400

n(Pb(OH)2) = m/M = 75 g/241.21 g/mol = 0.31 mol

m(NaOH) = n*M = 2*n(Pb(OH)2)*M(NaOH) = 24.8 g